Exercise 5-9 (Video) (Part Level Submission) The Sheridan Acres Inn is trying to determine its break-even point during its of(b) Determine the inns break-even point in dollars. Break-even point in
Answer #1

Variable costs=Maid service+Other costs

=(6+12)=$18 per room

Total fixed cost=Salaries+Utilities+Depreciation+Maintenance

=(4000+1500+1300+832)=$7632

Contribution margin=Revenue-Variable cost  

=36-18=$18 per room

Breakeven=Fixed expenses/Contribution margin

=7632/18

=424 rooms

=(424*36)

=$15264